OpenStudy (anonymous):

Solve the compound inequality. Express the solution using interval notation. n-4/3>=1 and 6n-3/3<=21

OpenStudy (anonymous):

\[n-\frac{4}{3}\geq1\] \[n\geq 1+\frac{4}{3}\] \[n\geq \frac{7}{3}\]

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

i bet this is \[\frac{n-4}{3}\geq 1\] yes?

OpenStudy (anonymous):

in which case \[n-4\geq 3\] \[n\geq7\]

OpenStudy (anonymous):

\[\frac{6n-3}{3}\leq21\] \[6n-3\leq 63\] \[6n\leq 66\] \[n\leq 11\]

OpenStudy (anonymous):

so you have \[7\leq n \leq 11\]

OpenStudy (anonymous):

interval notation is [7,11]